# Needed computations and computational capabilities for stellarators

### v1

@@ I Introduction
This paper points out areas in which computations could circumvent major issues in the development of fusion energy. Many of these issues were discussed in 2015 in the paper Stellarator design Boozer:design . For fusion energy to be rapidly and efficiently developed, a dynamic identification and study of such issues is needed.
Stellarators uniquely allow computational design to circumvent major issues. The coils that define the magnetic configuration make stellarator plasmas externally controlled to an extent unmatched by any other fusion concept, magnetic or inertial. The focus of stellarator design need not be extrapolation, with its conservatism and dangers from changing physics regimes, see Section 2 of Boozer:CO2-Stell , but innovative solutions.
@@
$\vec{B}=\frac{\mu_{0}G_{0}}{2\pi R}\hat{\varphi}+\vec{\nabla}\phi.$ (1)
$G_{0}$ is the number of Amperes of poloidal current, which comes up through the hole in the torus, and $\phi$ can be given by a single-valued solution to Laplace’s equation $\nabla^{2}\phi=0$. A Neumann boundary condition, $\hat{n}\cdot\vec{\nabla}\phi$, on any surface that encloses the toroidal plasma but excludes the coils, specifies $\varphi$ uniquely except for an irrelevant additive constant. Equivalently, $\phi$ can be represented by magnetic dipoles oriented normal to the enclosing surface. In the limit of a dense coverage of the enclosing surface by dipoles, the dipoles represent a single-valued current potential, Appendix A. Localized currents can be represented by patches of dipoles, which are a discrete representation of a localized current potential, called a current potential patch.
@@
Two features of optimized stellarators, which appear to be important for circumventing coil issues, have had little exploitation. The first is a result of Ku and Boozer in 2010 and 2011 Ku-Boozer:2010 ; Boozer-Ku:control2011 . Only a small fraction of the possible linear external perturbations $\delta\vec{B}\cdot\hat{n}$ at the surface of an optimized stellarator equilibria have an effect on the degree of physics optimization. The ones that do, such as the ones that affect global neoclassical transport, can be organized into external field distributions that decrease exponentially in importance from one distribution to the next. The second is a result of Todd Elder in his doctoral thesis Elder:thesis2023 under Allen Boozer and more completely in Elder-Boozer:2024 . The general field of Equation (1) with dipoles covering only $\sim 25\%$ of an enclosing surface at the location of the HSX stellarator coils can accurately reproduce their field on the plasma.
The primary focus of this paper is on unexploited freedom in the design of magnetic field coils. In particular, the focus is on curl-free models of stellarator plasmas, that could circumvent specific issues in fusion development. The primary reason for this focus is that innovative coil designs that produce curl-free plasma configurations can be determined with a reliability limited only by the competence of the group doing the calculation. These configurations can be perturbatively extended to full pressure models. There will be some discussion in the last sections of physics issues that affect stellarator design but have received little attention.
@@ II Fundamental coil choices
In 1972 Wobig and Rehker Modulars:1972 published a remarkable discovery in stellarator design. The effect of helical coils can be produced by modular coils. Modular coils, which are shaped toroidal field coils, can make both the toroidal and the helical fields required for a stellarator. Their much smaller size compared to helical coils simplifies the construction of stellarators. Modular coils revolutionized stellarator design.
@@
Nevertheless, modular coils have issues. They restrict plasma-chamber access, have intrinsic ripple, and provide an unnecessary restriction on plasma-coil separation. The plasma-coil separation issue can be understood in a straight-cylindrical $(r,\theta,z)$ model. The toroidal or $\hat{z}$-directed field and the number of Amperes required to produce it are independent of $r$. Fields that depend on the poloidal angle $\theta$, as in $\sin m\theta$, increase as $r^{m-1}$ as do the magnitude of currents required to produce them. Such fields are required in a stellarator: $m=2$ for ellipticity, $m=3$ for triangularity, etc. The implication is that modular coils for stellarators become more convoluted in shape the further back they are from the plasma. Current potential patches may allow a larger plasma-coil separation than can modular coils alone. The importance of this separation is discussed in Section IV, and the issue was studied in Todd Elder’s thesis Elder:thesis2023 .
@@
The independent use of a large number of coils to control plasma configurations has a fundamental mathematical limitation related to the condition number of matrices. This limitation is explained in Section V and can and should be studied computationally.
The magnetic fields associated with current-potential patches on an enclosing surface require coils be engineered. Attractive compact coils presumably exist when the magnetic field that a current-potential patch produces has a field strength that is highly constrained in its magnitude and its spatial gradients. Attractive designs for compact coil sets, Figure (1), that can be easily removed along with sections of the enclosing wall, whenever plasma chamber access is required, can and should be computationally designed.
@@
III Coils allowing easy plasma chamber access
There is no requirement that the plasma-encircling coils, such as modular coils, be chosen to ensure $\vec{B}\cdot\hat{n}=0$ on the plasma surface. But, they should be chosen to ensure good plasma access, a large separation from the plasma, Section IV, and may be chosen using other criteria such as simplicity. Construction difficulties associated with maintaining high spatial precision in the coils can be greatly reduced and will be discussed in Section VI.
[…]
Two concepts for open access to the plasma chamber are: (1) large spaces between coils and (2) non-plasma encircling coils that can be removed along with a large section of the chamber wall, Figure (1).
@@ III Coils allowing easy plasma chamber access
The use of non-plasma encircling coils mounted on removable wall sections have few fundamental constraints. Coils that can be removed along with the part of the vessel with which they are associated can be represented by current potential patches, which are patches of magnetic dipoles Elder:thesis2023 ; Elder-Boozer:2024 . When a section of the vessel is removed for plasma-chamber access, it must be placed back into position with a high accuracy to obtain a vacuum seal—probably with more accuracy than is required for the placement of the coils.
@@
IV Coils with maximal separation from the plasma
@@
The second is a measure of how efficiently distant coils can support a magnetic surface of that shape. For curl-free magnetic configurations, external coils must produce a field normal to the reference surface $b(\theta,\varphi)=-(\mu_{0}G_{0}/2\pi R)\hat{\varphi}\cdot\hat{n}$, where $\hat{n}$ is the unit normal to the reference surface. Coil efficiency drops exponentially with separation Boozer:RMP . The coefficient in the exponential is approximately given by the wavenumber of the variation of $b$ in the reference surface Landreman:field-gradient . Even crude measures, such as the strength of the dipoles required to produce the field, $\vec{d}^{\dagger}\cdot\vec{d}$ of Appendix A, provide a useful measure.
@@
V Coils allowing flexibility
There are two types of external perturbations: (1) ones that are consistent with the periodicity of the optimized configuration and (2) those that are not. The first type are considered changes in the configuration while the second type are considered to be error fields, the effect of construction inaccuracies. The orders of magnitude differences in configuration sensitivity to changes in different external magnetic field distributions should be recognized in stellarator coil design. Designing coils by minimizing $\vec{B}\cdot\hat{n}$ on a given surface ignores these differences.
@@ V Coils allowing flexibility
Unfortunately, there is a mathematical limit on the number of currents that can be adjusted with sufficient accuracy to control physics properties. This limit is approximately given by the condition number of a matrix.
@@
To understand the condition-number problem, define a matrix vector $\vec{P}$ that consists of numbers that define plasma properties—for example, the shape parameters of the plasma surface. $\vec{P}$ is related to the matrix vector that gives the control currents $\vec{I}$ by a control matrix, $\vec{P}=\tensor{C}\cdot\vec{I}$. The required current to obtain a specified set of properties is $\vec{I}=\tensor{C}^{-1}\cdot\vec{P}$. An important question is when a small change is made in $\vec{P}$ how large is the maximum change in $\vec{I}$ that may be required. There is a fundamental control problem when the relative change in the current $||\delta\vec{I}||/||\vec{I}||$ is enormous when compared to the desired change in the plasma properties $||\delta\vec{P}||/||\vec{P}||$.
The maximum possible fractional change in the currents relative to a fractional change in the desired properties is
@@
$\hskip 36.135pt=Max\left\{||\tensor{C}^{-1}||\right\}Max\left\{||\tensor{C}||\right\}.$ (2)
@@
The problem arises when one wishes to control a component of $\vec{P}$ that is associated with a small singular value of $\tensor{C}$, which means a component of $\vec{I}$ must be very large to drive it. Unless that component of $\vec{I}$ can be made orthogonal to all components of $\vec{I}$ associated with large singular values, one cannot control the component of $\vec{P}$ that one wants without making a large unwanted change in a different component of $\vec{P}$. The hypersensitivity of the currents in the coils relative to the plasma parameters that are to be controlled can be reduced by eliminating the difficult components in $\delta\vec{P}$.
The actual sensitivity in plasma control can be much smaller, than the condition-number estimate. Computer simulations can estimate what types of control are practical and how to design the control coils to minimize their sensitivity to important changes in $\vec{P}$.
VI Coils that reduce precision requirements
The calculation of control or error field sensitivities has been largely ignored in stellarator-design studies as well as in tokamak studies. Nevertheless, it is difficult to grasp how a coil system can be considered to be optimized without their inclusion.
Once an attractive stellarator equilibrium is found, one should determine what $\delta\vec{B}\cdot\hat{n}$’s greatly degrade the attractiveness of that equilibrium. These $\delta\vec{B}\cdot\hat{n}$’s, which frequently have an $n=1$ toroidal Fourier dependence, are the error fields that should be made controllable. They define a matrix vector $\vec{b}_{e}$ of error fields, which can be controlled by current potentials defined by $\tensor{M}_{c}\cdot\vec{d}=\vec{b}_{e}$, Appendix A. A Singular Value Decomposition of $\tensor{M}_{c}$ allows one to write $\vec{d}_{c}\equiv\tensor{V}_{c}^{\dagger}\cdot\vec{d}$, $\vec{b}_{c}\equiv\tensor{U}_{c}^{\dagger}\cdot\vec{b}_{e}$, with $\vec{d}_{c}=\tensor{S}_{c}^{-1}\cdot\vec{b}_{c}$, where $\tensor{S}_{c}$ is the diagonal matrix of the singular values.
When $k_{c}$ error fields are to be controlled, current potential patches must be found that can independently control the error fields. This becomes increasingly impractical, Section V, the larger the condition number, the ratio of the largest to the smallest non-zero singular value in $\tensor{S}_{c}$.
Error field control becomes impractical when the expected errors in the determination of the equilibrium are sufficiently large that the $k_{c}$ error fields $\vec{b}_{e}$ cannot be reliably distinguished by the error-field current-potential patches. It is this limit that determines an accuracy with which stellarators must be built. This may well be a lower accuracy requirement than that for minimal-cost mechanical construction—bolts must fit through holes and a vacuum must be achievable.
VII Coil designs consistent with major configuration changes
As discussed in Section V, plasma sensitivity to small changes in the external magnetic field defines external field distributions that should be controllable in order to have experimental flexibility. However, far more radical changes of the plasma configuration are possible—for example changing the plasma shape from a quasi-isodynamic to a quasi-helically symmetric equilibrium. A fundamental limit on radical changes is set by the plasma not intercepting non-replaceable parts of the chamber walls. More subtle limits are set by the locations of divertors and diagnostics.
Todd Elder’s result that approximately 25% coverage of an enclosing surface by current potential patches allows the accurate definition of a quasi-helically symmetric equilibrium suggests two strategies: (1) Define sufficient current-potential patches that by changing which are driven a major change in configuration can be made. (2) When the current-potential patches are on removable sections of the chamber walls, one set of wall inserts carrying a current-potential patch could be replaced by another. The first option sounds preferable, but changes in the surrounding walls may be sufficiently great that the second option is required.
VIII Configuration possibilities
@@
VIII.1 Possible divertor designs
One such goal in the design and control of plasma divertors. An important computational study would be what possibilities these remaining degrees of freedom allow in divertor magnetic configurations. In particular, it is important to know which field lines that enter the plasma chamber and make a large number of toroidal transits before exiting come close to the region of magnetic surfaces that define the plasma. These are the lines that determine the divertor properties of that configuration.
There are two basic types of stellarator divertors: (1) the island divertor as on W7-X W7-X:divertor and (2) the non-resonant divertor.
Non-resonant divertors appear naturally in optimized stellarators due to locations of sharp curvature of the magnetic surfaces near the plasma edge, Bader et al Bader:2018 . Due to $\vec{\nabla}\cdot\vec{B}=0$, these are also the locations at which the separation between neighboring magnetic surfaces becomes large. Boozer and Punjabi Boozer-Punjabi:2018 used these features to derive an analytic magnetic field line Hamiltonian, the poloidal magnetic flux as a function to the toroidal magnetic flux and two angles, $\psi_{p}(\psi,\theta,\varphi)$. They have used this Hamiltonian in studies of topological features of non-resonant divertors, such as cantori, turnstiles, and magnetic field line chaos.

### v2

@@ I Introduction
Stellarators uniquely allow computational design to circumvent major issues. Many were discussed in 2015 in the paper Stellarator design Boozer:design . The coils that define the magnetic configuration make stellarator plasmas externally controlled to an extent unmatched by any other fusion concept, magnetic or inertial. The focus of stellarator design need not be extrapolation, with its conservatism and dangers from changing physics regimes, see Section 2 of Boozer:CO2-Stell , but innovative solutions, which could speed the development of attractive fusion power.
@@
$\vec{B}_{x}=\frac{\mu_{0}G_{0}}{2\pi R}\hat{\varphi}+\vec{\nabla}\phi.$ (1)
$G_{0}$ is the number of Amperes of poloidal current, which comes up through the hole in the torus, and $\phi$ can be given by a single-valued solution to Laplace’s equation $\nabla^{2}\phi=0$. A Neumann boundary condition, $\hat{n}\cdot\vec{\nabla}\phi$, on any surface that encloses the toroidal plasma but excludes the coils, specifies $\phi$ uniquely except for an irrelevant additive constant. Equivalently, $\phi$ can be represented by magnetic dipoles oriented normal to the enclosing surface. In the limit of a dense coverage of the enclosing surface by dipoles, the dipoles represent a single-valued current potential, Appendix A. Localized currents can be represented by patches of dipoles, which are a discrete representation of a localized current potential, called a current potential patch by Todd Elder Elder-Boozer:2024 .
@@
Two features of optimized stellarators, which appear to be important for circumventing coil issues, have had little exploitation. The first is a result of Ku and Boozer in 2010 and 2011 Ku-Boozer:2010 ; Boozer-Ku:control2011 . Only a small fraction of the possible linear external perturbations $\delta\vec{B}_{x}\cdot\hat{n}$ at the surface of an optimized stellarator equilibria have an effect on the degree of physics optimization. The ones that do, such as the ones that affect global neoclassical transport, can be organized into external field distributions that decrease exponentially in importance from one distribution to the next. The second is a result of Elder and Boozer Elder-Boozer:2024 . The general field of Equation (1) with dipoles covering only $\sim 25\%$ of an enclosing surface at the location of the HSX stellarator coils can accurately reproduce their field on the plasma.
@@ II Fundamental coil choices
In 1972 Wobig and Rehker Modulars:1972 published a discovery that revolutionized stellarator coil design. The effect of helical coils can be produced by modular coils. Modular coils, which are shaped toroidal field coils, can make both the toroidal and the helical fields required for a stellarator. Their much smaller size compared to helical coils simplifies the construction of stellarators.
@@
Nevertheless, modular coils have issues. They restrict plasma-chamber access, have intrinsic ripple, and provide an unnecessary restriction on plasma-coil separation. The plasma-coil separation issue can be understood in a straight-cylindrical $(r,\theta,z)$ model. The toroidal, or $\hat{z}$-directed, field and the number of Amperes required to produce it are independent of $r$. Fields that depend on the poloidal angle $\theta$, as in $\sin m\theta$, increase as $r^{m-1}$ as do the magnitude of currents required to produce them. Such fields are required in a stellarator: $m=2$ for ellipticity, $m=3$ for triangularity, etc. The implication is that modular coils for stellarators become more convoluted in shape the further back they are from the plasma. Current potential patches may allow a larger plasma-coil separation than can modular coils alone. The importance of this separation is discussed in Section II.2, and the issue was studied in Todd Elder’s thesis Elder:thesis2023 .
@@
The independent use of a large number of coils to control plasma configurations has a fundamental mathematical limitation related to the condition number of matrices. This limitation is explained in Section III and can and should be studied computationally.
The magnetic fields associated with current-potential patches on an enclosing surface require coils be engineered. Attractive compact coils presumably exist when the magnetic field that a current-potential patch produces has a field strength that is highly constrained in its magnitude and its spatial gradients.
II.1 Coils allowing easy plasma chamber access
There is no requirement that the plasma-encircling coils, such as modular coils, be chosen to ensure $\vec{B}\cdot\hat{n}=0$ on the plasma surface. But, they should be chosen to ensure good plasma access, a large separation from the plasma, Section II.2, and may be chosen using other criteria such as simplicity. Construction difficulties associated with maintaining high spatial precision in the coils can be greatly reduced and will be discussed in Section II.4.
Access affects both the cost of fusion power and the speed with which fusion systems can be developed.
The 14 MeV neutrons from deuterium-tritium (DT) fusion reactions degrade materials unacceptably after a certain fluence, which is approximately 10 MW yr/m2. Fusion energy is cheaper the greater the power density, MW/m2 through the walls until material limits are reached. A major limitation on the fraction of time a fusion power plant can actually produce power, its duty factor, is the speed with which neutron-degraded plasma-chamber components can be replaced. Even in ITER ITER hot Cell: 2017 , the replacement of degraded components is complicated by both their high level of radioactivity and by the requirement of maintaining confinement of tritium. The required replacement time is minimized by open plasma-chamber access.
@@ II.1 Coils allowing easy plasma chamber access
Two concepts for open access to the plasma chamber are: (1) large spaces between coils and (2) non-plasma encircling coils that can be removed along with a large section of the chamber wall Boozer:2009 , Figure (1).
@@
The use of non-plasma encircling coils mounted on removable wall sections have few fundamental constraints. Coils that can be removed along with the part of the vessel with which they are associated can be represented by current potential patches, which are patches of magnetic dipoles Elder-Boozer:2024 . When a section of the vessel is removed for plasma-chamber access, it must be placed back into position with a high accuracy to obtain a vacuum seal—probably with more accuracy than is required for the placement of the coils.
@@
II.2 Coils with maximal separation from the plasma
@@
The second is a measure of how efficiently distant coils can support a magnetic surface of that shape. For curl-free magnetic configurations, external coils must produce a field normal to the reference surface $b(\theta,\varphi)=-(\mu_{0}G_{0}/2\pi R)\hat{\varphi}\cdot\hat{n}$, where $\hat{n}$ is the unit normal to the reference surface. Coil efficiency drops exponentially with separation Boozer:RMP . The coefficient in the exponential is approximately given by the wavenumber of the variation of $b$ in the reference surface Landreman:field-gradient . This result follows from a two-dimensional model; $\nabla^{2}\phi=0$ implies $\phi$ has the typical dependence $\sin(k_{y}y)\exp(-k_{x}x)$ with $k_{x}=k_{y}$. The exponential drop in efficiency implies that crude measures, such as the strength of the dipoles required to produce the field, $\vec{d}^{\dagger}\cdot\vec{d}$ of Appendix A, provide a useful measure.
@@
II.3 Coils allowing flexibility
There are two types of external perturbations: (1) ones that are consistent with the periodicity of the optimized configuration and (2) those that are not. The first type are considered changes in the configuration while the second type are considered to be error fields, the effect of construction inaccuracies. The orders of magnitude differences in configuration sensitivity to changes in different external magnetic field distributions Ku-Boozer:2010 ; Boozer-Ku:control2011 should be recognized in stellarator coil design. Designing coils by minimizing $\vec{B}\cdot\hat{n}$ on a given surface ignores these differences.
@@ II.3 Coils allowing flexibility
Unfortunately, there is a mathematical limit on the number of currents that can be adjusted with sufficient accuracy to control physics properties. This limit is approximately given by the condition number of a matrix, Section III.
II.4 Coils that reduce precision requirements
The calculation of control or error field sensitivities has been largely ignored in stellarator-design studies as well as in tokamak studies. Nevertheless, it is difficult to grasp how a coil system can be considered to be optimized without such a study.
[…]
There are two strategies for reducing the precision requirements of coil construction. First, the plasma optimization can be used to reduce the plasma sensitivity to the worst magnetic-field error, which is defined by $\kappa_{1}$. This efficiently accomplishes what is accomplished by a stochastic optimization Stochastic optimization .
Secondly, error-field correction coils can be installed that can independently drive the various magnetic field distributions to which the plasma is sensitive. These coils can make the total external field in distributions to which the plasma is sensitive zero. External field errors are not cancelled. The external drive just makes the externally-produced magnetic field distributions to which the plasma is sensitive zero. Field errors to which the plasma is not sensitive are driven in this process, so there is maximum number of error field distributions that can be controlled in this way. Fortunately, the plasma sensitivity appears to drop exponentially, so only a few error field distributions need to be controlled to greatly reduce the precision requirements on the coils.
II.5 Coil designs consistent with major configuration changes
As discussed in Section II.3, plasma sensitivity to small changes in the external magnetic field defines external field distributions that should be controllable in order to have experimental flexibility. However, far more radical changes of the plasma configuration are possible—for example changing the plasma shape from a quasi-isodynamic to a quasi-helically symmetric equilibrium. A fundamental limit on radical changes is set by the plasma not intercepting non-replaceable parts of the chamber walls. More subtle limits are set by the locations of divertors and diagnostics.
Todd Elder’s result that approximately 25% coverage of an enclosing surface by current potential patches allows the accurate definition of a quasi-helically symmetric equilibrium suggests two strategies: (1) Define current-potential patches that are consistent with major changes in the configuration. (2) When the current-potential patches are on removable sections of the chamber walls, one set of wall inserts carrying a current-potential patch could be replaced by another. The first option sounds preferable, but changes in the surrounding walls may be sufficiently great that the second option is required.
III Control Matrix Issues
The control matrix $\tensor{C}$ relates a matrix vector $\vec{I}$ that has elements equal to the currents in the control coils to the parameters that one one wishes to control, the important perturbations to the external magnetic field $\delta\vec{B}_{x}\cdot\hat{n}$ on the plasma surface. That is
$\vec{P}=\tensor{C}\cdot\vec{I}.$ (2)
The external field perturbations on the plasma surface are of two types: (1) ones that are consistent with the periodicity of the optimized configuration, Section II.3, and (2) those that are not—the error fields, Section II.4.
The mathematics of the two types of perturbations is somewhat different. The first type is defined by external magnetic fields $\delta\vec{B}_{x}\cdot\hat{n}$ associated with large terms in the Hessian matrix. The second type is defined by the current potentials $\kappa_{1}$, $\kappa_{2},\cdots$ to which the plasma is sensitive. For both types, the current potentials that would produce the precise perturbation would be be non-zero over a large fraction of the surface on which they are defined and are generally sensitive to the plasma state that one is trying to control.
To obtain practical coils, only patches of current potential can be retained, which should be in the places where the current potentials required for control are largest. Although patches of current potential can control the quantities that are to be controlled, they do not allow the independent control of all possible $\delta\vec{B}_{x}\cdot\hat{n}$ distributions. Compact current-potential patches produce crosstalk between the drives of the various $\delta\vec{B}_{x}\cdot\hat{n}$ distributions.
There is a subtle mathematical problem. How is the choice of control coils to be made in order to have a control matrix $\tensor{C}$ that optimizes the control of the most important $\delta\vec{B}_{x}\cdot\hat{n}$ distributions over the broad range of plasma states that an experiment is designed to access?
@@ III Control Matrix Issues
An important question is when a small change is to be made in $\vec{P}$ how large is the maximum change in $\vec{I}$ that may be required. There is a fundamental control problem when the relative change in the current $||\delta\vec{I}||/||\vec{I}||$ is enormous when compared to the desired change in the properties $||\delta\vec{P}||/||\vec{P}||$. The maximum possible fractional change in the currents relative to a fractional change in the desired properties is
@@
$\hskip 36.135pt\leq Max\left\{||\tensor{C}^{-1}||\right\}Max\left\{||\tensor{C}||\right\}.$ (3)
@@
A problem arises when one wishes to control a component of $\vec{P}$ that is associated with a small singular value of $\tensor{C}$, which means a component of $\vec{I}$ must be very large to drive it. Unless that component of $\vec{I}$ can be made orthogonal to all components of $\vec{I}$ associated with large singular values of $\tensor{C}$, one cannot control the component of $\vec{P}$ that one wants without making a large unwanted change in a different component of $\vec{P}$. The hypersensitivity of the currents in the coils relative to the plasma parameters that are to be controlled can be reduced by limiting the components in $\vec{P}$ that one is trying to control to just the most important. The results of Ku and Boozer Ku-Boozer:2010 ; Boozer-Ku:control2011 imply that only a few distributions of $\vec{B}_{x}\cdot\hat{n}$ require careful control.
A clever choice of the control coils can make the actual sensitivity in plasma control smaller, even much smaller, than the condition-number estimate. Computer simulations can estimate what types of control are practical and how to design the control coils to provide maximal independent control of the most important elements in $\vec{P}$ over a variety of plasma states. The limitations and the possibilities for control are clarified by the discussion in Appendix A.3 on the specification of external fields.
IV Configuration possibilities
@@
IV.1 Possible divertor designs
One such goal in the design and control of plasma divertors. An important computational study would be what possibilities these remaining degrees of freedom allow for divertors.
There are two basic types of stellarator divertors: (1) the island divertor as on W7-X W7-X:divertor and (2) the non-resonant divertor. The presence of a strong bootstrap current is thought to make island divertors, the type of divertor used on W-7X, impractical. This may be less true in a power plant than in an experimental facility. A quasi-steady-state power plant could operate in a single plasma state with a known edge transform. Both the edge transform and surface shape can be controlled by external coils during startup. The rotational transform can be controlled by poloidal field coils forming a transformer that allows a controlled change in poloidal flux in the central hole of the torus.